Assuming that the price deceases from $10 to $5 and the quantity demanded increases from 20 units to 30 units, which of the following is true


the demand is inelastic


the price elasticity of demand is 1


demand is unit elastic


the price is perfectly elastic


the price elasticity of demand is -1




1
Expert's answer
2021-01-07T07:58:56-0500

If the price deceases from $10 to $5 and the quantity demanded increases from 20 units to 30 units, then the elasticity of demand is:

"Ed = \frac {30-20} {5-10} × \frac {5+10} {30+20} = -0.6,"

so the correct answer is "the demand is inelastic".